Circular No. 8354

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V1647 ORIONIS = IRAS 05436–0007 N. N. Samus, Institute of Astronomy, Moscow, informs us that the

designation V1647 Ori has been given to the variable star associated with IRAS 05436–0007 (cf. IAUC 8284, 8289, 8290, 8303, 8306).

S. Vig, S. K. Ghosh, and D. K. Ojha, Tata Institute of Fundamental Research, Mumbai, in collaboration with V. K. Kulkarni (National Centre of Radio Astrophysics, Pune), report that a radio-continuum observation of IRAS 05436-0007, obtained at 1272 MHz using the Giant Metrewave Radio Telescope on Feb. 14.5 UT, shows no radio emission within a 1′.5 × 1′.5 region around IRAS 05436–0007 (α = 5^{h}46^{m}13^{s}.1, δ = −0^{o}06′05′′, equinox 2000.0; from 2MASS), with a 5σ upper limit of 0.15 mJy/beam (synthesized beam size 5′′.61 × 2′′.72, p.a. −48^{o}.91). All of the sources from the NRAO/VLA Sky Survey Catalog within this region (covering a 25′- diameter circular area centered on the nebula) were detected.

(4492) DEBUSSY Further to IAUC 8265 and 8292, R. Behrend writes (on behalf of four-

teen others) that photometric observations of (4492) obtained during 2002 Oct. 30–Dec. 4 and 2004 Mar. 18–May 18 show a regularly shaped lightcurve with period 1.108 day and amplitude 0.5 mag, with sharp attenuations characteristic of mutual eclipse/occultation events observed at minima of the regular lightcurve having depth ∼ 0.6 mag and duration ∼ 2.5 hr — suggesting that this is probably a binary system, with properties similar to those of (1089) and (1313).

COMETS C/2003 V6 AND C/2003 V7 (SOHO) Further to IAUC 8351, K. Battams reports measurements for addi-

tional Kreutz sungrazing comets found on SOHO website C2 images by T. Hoffman (C/2003 V6) and R. Matson (C/2003 V7). The reductions by R. Kisala and orbital elements by B. G. Marsden appear on MPEC 2004-K34.

Comet  2003 UT  α_{2000}  δ_{2000}
C/2003 V6 Nov. 9.564  14^{h}53.^{m}4  −18^{o}10′
C/2003 V7  11.896  15 03.5  −18 57

SUPERNOVA 2004ch IN NGC 5612 Corrigendum. On IAUC 8353, line 2, for May 12.619 read May 28.63

2004 June 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green